A block sits on the floor. (a) What is the magnitude of the frictional force on it from the floor? (b) If a horizontal force of 5 N is now applied to the block, but it does not move, what is the magnitude of the frictional force on it? (c) If the maximum value fs,max of the static frictional force on the block is 10 N, and you apply a 12 N force, what is the magnitude of the frictional force, and how much will a 500-g mass will accelerate?
Answer #1

(a)

As the block sits on the floor and there no force applied to it. So there is no frictional force to oppose the motion.

Answer : Zero

(b)

Now , 5N horizontal force is applied to the block but the block does not move. It means there is a frictional force equal to the horizontal force (5 N) which is acting opposite to the direction of applied force.

Answer : 5 N

(c)

Frictional force (Fs) = 10 N

Applied force (F)= 12 N

Net force (Fnet) = F - Fs = 12 N - 10 N = 2 N

According to the newton's second law of motion , we have :

Fnet = Ma

2 N = 0.5 kg × a

i.e. a = 2 N / 0.5 kg = 4 m/s2

Answer : Frictional force = 10 N ; acceleration= 4 m/s2